A reaction happens with 75 grams of acetic acid (C2O2H4) and an unknown metal, resulting in hydrogen gas and a solid metal with a mass of 113.4 g. What is the unknown metal?


1
Expert's answer
2021-01-07T05:31:24-0500

Assuming M is a reactive univalent metal, its reaction with acetic acid is;

"2CH_3COOH_{(aq)} + 2M_{(s)} \to 2CH_3COOM_{(aq)} + H_{2(g)}"


According to this reaction,

2 moles of acetic acid reacts with 2 moles of the metal


1 mole of acetic acid = 60g

x mole of acetic acid = 75/60 = 1.25 mol


Therefore, 1.25 moles of acetic acid reacts with 1.25 moles of the metal acetate


1.25 moles of the metal = 113.4g

Therefore, 1 mole of the metal = 90.72g ("\dfrac{113.4×1}{1.25}")


This means the molar mass of the metal is 90.72g/mol


Upon consultation of the periodic table, the metallic element with a molar mass of approximately 91g/mol is Zirconium with the symbol Zr and an atomic number of 40.